OpenStudy (anonymous):

What is the equation of the line passing through (–3, –9) and parallel to the line y = x + 6 in standard form?

OpenStudy (anonymous):

we know parrallel lines have the same slope, right?

OpenStudy (anonymous):

the slopes are equal the slope of y=x+6 is m=1. therefore the slope of parallel line is also 1

OpenStudy (anonymous):

The equation of requested line is: y + 9 = 1*(x+3) y + 9 = x + 3 y = x - 6

OpenStudy (anonymous):

x-y-6=0